# Weak $\mathbb{Z}_{2}$ Supertopology

### v1

@@ Abstract
Crystal symmetries can enforce all bands of a material to be topological, a property that is commonly referred to as “supertopology”. Here, we determine the symmetry-enforced $\mathbb{Z}_{2}$ supertopologies of non-magnetic centrosymmetric materials with weak and strong spin-orbit coupling (SOC). For weak (i.e., negligible) SOC, crystal symmetries can enforce Dirac nodal lines protected by a $\pi$-Berry phase, while for strong SOC, crystal symmetries can give rise to nontrival weak $\mathbb{Z}_{2}$ topologies in 2D subplanes of the 3D Brillouin zone. We catalogue all centrosymmetric space groups whose symmetries enforce these $\mathbb{Z}_{2}$ supertopologies. Suitable material realizations are identified and experimental signatures of the supertopologies, such as surfaces states and quantum Hall responses, are being discussed.
@@ I Introduction
In the last decade and a half, the study of materials with non-trivial band topology has become an important field of physics Wieder et al. (2022); Hasan and Kane (2010); Chiu et al. (2016). The exotic properties of these materials, such as chiral band crossings, anomalous surface states, and topological transport, are described by topological invariants, e.g., Chern numbers or $\mathbb{Z}_{2}$ indices Kane and Mele (2005). In order to understand the origin of these topological band features, it is necessary to investigate the connection between symmetry and topology: The crystal symmetries do not only lead to symmetry-related copies of the topological band features, but they enforce by themselves certain nontrivial band topologies. For example, screw rotations enforce Weyl points on rotation axes Hirschmann et al. (2021); Leonhardt et al. (2021); Alpin et al. (2023), while glide mirrors enforce nodal lines Chan et al. (2019); Fang et al. (2016); Bian et al. (2016); Xie et al. (2021); Takahashi et al. (2017). This property is referred to as “symmetry-enforced topology” Watanabe et al. (2016); Takahashi et al. (2017); Zhao and Schnyder (2016); Young and Wieder (2017); Chan et al. (2019); Hirschmann et al. (2021); Leonhardt et al. (2021); Wilde et al. (2021); Micklitz and Norman (2017); Ono et al. (2021); Jiang et al. (2024) and guarantees that bands must be topological irrespective of the chemical composition and other details of the material. The principle of symmetry-enforced topology can be used to design or search for new materials with the desired topological characteristics Leonhardt et al. (2021). Another intriguing class of materials is called “supertopological materials” Vergniory et al. (2022), for which all connected bands isolated in energy have a stable topological invariant. To obtain topological responses in supertopological materials, a precise tuning of the electronic filling is not necessary, since all bands are topologically nontrivial.
@@ III.1 Space Group 15 (C2/c)
We constructed a tight-binding (TB) model with symmetries of the space group 15 with and without SOC Fig. 1 (b)-(c) [see Appendix E.1]. There is an enforced crossing inside the plane, corresponding to the piercing nodal line. The degeneracy is split under SOC opening a gap inside the L-V plane. For the constructed TB models, we introduced a termination direction along the $y$-axis to observe surface states due to non-trivial topology, see Fig. 1 (b)-(c).
@@ IV.1 Space Group 135 (P42/mbc)
We discuss the approach considering the space group 135. The spinless representations of glide mirror symmetries are given by
@@ VI Conclusions
In this work, we investigated the origin of supertopology. We found 17 SGs enforcing topological invariants for every band independently of material characteristics. In several space groups, the symmetry-enforced Dirac nodal line (and its surface state) is the parent state for the quantum spin Hall state if SOC is added.
@@
In the description of the enforced Kane-Mele invariant, we assumed the presence of the bulk band gap in a plane. This assumption does not demand the bulk gap everywhere in the full BZ. This implies the possible coexistence of the bulk Fermi surface with boundary resonances Bergman and Refael (2010). If the origin of the Fermi surface is the symmetry-enforced connectivity between the bands Vergniory et al. (2022), the bulk gap can be opened with the breaking of the crystalline symmetries. Unless there is a band inversion in the non-trivial plane, the invariant must be preserved and non-trivial. This can be an approach to create symmetry-enforced weak topological insulators.
Approaches used in the paper can be applied to study representation-enforced topology Parshukov . In this case, one can consider a particular Wyckoff position and assume that bands induced from each of the site symmetry groups do not mix. With these conditions the classification of the space groups can be extended to the classification of the Wyckoff positions that induce bands carrying weak and strong $\mathbb{Z}_{2}$, and higher-order $\mathbb{Z}_{4}$ topologies Khalaf et al. (2018); Schindler et al. (2018); Benalcazar et al. (2017); Parshukov .
The discussed symmetry indicators can be applied to magnetic systems with $P$ and anti-unitary symmetries. The anti-unitary symmetry (time-reversal with a crystalline operation) must quantize $\mathbb{Z}_{2}$ topological invariant Mong et al. (2010), and the inversion eigenvalues indicate a non-trivial phase. Thus, the material search can be extended to magnetic phases.
With the addition of chiral symmetry, 2D electron systems can host higher-order topology Langbehn et al. (2017). Therefore, symmetry-enforced higher-order topological phases are possible and require further investigation.
We study the 2D symmetry indicators as the 3D counterparts can not be enforced due to material dependence of the inversion eigenvalues at $\Gamma$ point. However, in higher dimensional systems, the approach can be applied to enforce 3D symmetry indicators.

### v2

@@ Abstract
Crystal symmetries can enforce all bands of a material to be topological, a property that is commonly referred to as “supertopology”. Here, we determine the symmetry-enforced $\mathbb{Z}_{2}$ supertopologies of non-magnetic centrosymmetric materials with weak and strong spin-orbit coupling (SOC). For weak (i.e., negligible) SOC, crystal symmetries can enforce Dirac nodal lines protected by a $\pi$-Berry phase, while for strong SOC, crystal symmetries can give rise to nontrival weak $\mathbb{Z}_{2}$ topologies in 2D subplanes of the 3D Brillouin zone. We catalogue all centrosymmetric space groups whose symmetries enforce these $\mathbb{Z}_{2}$ supertopologies. Suitable material realizations are identified and experimental signatures of the supertopologies, such as quantum spin Hall states, are being discussed.
@@ I Introduction
In the last decade and a half, the study of materials with non-trivial band topology has become an important field of physics Wieder et al. (2022); Hasan and Kane (2010); Chiu et al. (2016). The exotic properties of these materials, such as chiral band crossings, anomalous surface states, and topological transport, are described by topological invariants, e.g., Chern numbers or $\mathbb{Z}_{2}$ indices Kane and Mele (2005). In order to understand the origin of these topological band features, it is necessary to investigate the connection between symmetry and topology: The crystal symmetries do not only lead to symmetry-related copies of the topological band features, but they enforce by themselves certain nontrivial band topologies. For example, screw rotations enforce Weyl points on rotation axes Hirschmann et al. (2021); Leonhardt et al. (2021); Alpin et al. (2023), while glide mirrors enforce nodal lines Chan et al. (2019); Fang et al. (2016); Bian et al. (2016); Xie et al. (2021); Takahashi et al. (2017). This property is referred to as “symmetry-enforced topology” Watanabe et al. (2016); Takahashi et al. (2017); Zhao and Schnyder (2016); Young and Wieder (2017); Chan et al. (2019); Hirschmann et al. (2021); Leonhardt et al. (2021); Wilde et al. (2021); Micklitz and Norman (2017); Ono et al. (2021) and guarantees that bands must be topological irrespective of the chemical composition and other details of the material. The principle of symmetry-enforced topology can be used to design or search for new materials with the desired topological characteristics Leonhardt et al. (2021). Another intriguing class of materials is called “supertopological materials” Vergniory et al. (2022), for which all connected bands isolated in energy have a stable topological invariant. Similar property was studied in electric Jiang et al. (2024) and superconducting Fatemi et al. (2021); Herrig and Riwar (2022); Peyruchat et al. (2024) circuits, where all bands carry a non-trivial invariant. To obtain topological responses in supertopological materials, a precise tuning of the electronic filling is not necessary, since all bands are topologically nontrivial.
@@ III.1 Space Group 15 (C2/c)
We constructed a tight-binding (TB) model with symmetries of the space group 15 with and without SOC Fig. 1 (b)-(c) [see Appendix E.1]. There is an enforced crossing inside the plane, corresponding to the piercing nodal line. The degeneracy is split under SOC opening a gap inside the L-V plane. For the constructed TB models, we introduced a termination direction along the $y$-axis to observe surface states due to non-trivial topology, see Fig. 1 (b)-(c). The drumhead surface state bounded by the Dirac nodal line transforms under SOC into a Dirac cone edge state at the projected $\overline{\text{L}}$ point.
@@ IV.1 Space Group 135 (P42/mbc)
We discuss the approach considering the space group 135. The SG is shown to enforce a double Dirac point at the A point Wieder et al. (2016). Here we are interested in the M, X, R and Z points. The spinless representations of glide mirror symmetries are given by
@@ VI Conclusions
In our work, we investigated the conditions under which a material is supertopological, meaning that all its bands have a nontrivial topology. We focused on non-magnetic centrosymmetric materials, with both weak and strong spin-orbit coupling (SOC), and determined the symmetry conditions that enforce a supertopological band structure. Specifically, we identified and catalogued the centrosymmetric space groups whose symmetries enforce $\mathbb{Z}_{2}$ supertopologies. We found 17 space groups enforcing topological invariants for every band, independently of material characteristics (see Tables 1 and 2).
For materials with weak SOC, we demonstrated that the $\mathbb{Z}_{2}$ supertopology gives rise to Dirac nodal lines protected by a $\pi$-Berry phase. In contrast, for materials with strong SOC, the supertopology manifests itself by weak $\mathbb{Z}_{2}$ invariants in 2D subplanes of the 3D Brillouin zone with an associated quantum-spin Hall edge state. In several space groups, the symmetry-enforced Dirac nodal line (and its drumhead surface state) is the parent precursor state of the weak $\mathbb{Z}_{2}$ invariant (and the quantum-spin Hall edge state) that arises once SOC is turned on.
@@
In the description of the enforced $\mathbb{Z}_{2}$ invariant, we assumed the presence of a bulk band gap in a 2D plane. This assumption does not require a bulk gap everywhere in the full BZ. This implies the possible coexistence of a bulk Fermi surface with boundary resonances of the drumhead or quantum-spin Hall surface state Bergman and Refael (2010). We note that several of the found space groups are examples of “filling-enforced” gapless states Wieder and Kane (2016), like SG 135, which has a double Dirac point at the BZ corner Wieder et al. (2016). If the origin of the Fermi surface is the symmetry-enforced connectivity between the bands Vergniory et al. (2022), the bulk gap can be opened with the breaking of the crystalline symmetries. Unless there is a band inversion in the non-trivial plane, the invariant must be preserved and non-trivial. This can be an approach to create symmetry-enforced weak topological insulators.
The approach developed in this work can be extended, in a straightforward manner, to determine representation-enforced supertopologies Parshukov . That is, one can consider a particular Wyckoff position and assume that bands with their irreps induced from each of the site symmetry groups do not mix. In this way, one can classify Wyckoff positions and site symmetry groups that enforce weak and strong $\mathbb{Z}_{2}$, and higher-order $\mathbb{Z}_{4}$ topologies Khalaf et al. (2018); Schindler et al. (2018); Benalcazar et al. (2017); Parshukov . Furthermore, our approach can be applied also to magnetic systems with $P$ and anti-unitary symmetries. That is, also in magnetic materials anti-unitary symmetries (time-reversal with a crystalline operation) can quantize $\mathbb{Z}_{2}$ topological invariants Mong et al. (2010), in which case the inversion eigenvalues indicate the non-trivial phase. Performing such an analysis for all centrosymmetric magnetic space groups will give search criteria for novel magnetic materials with $\mathbb{Z}_{2}$ supertopologies.
Another avenue for future research is to add chiral (sublattice) symmetry to the game, whose presence can lead to 2D electron systems with higher-order topology Langbehn et al. (2017). It would be interesting to figure out the symmetry conditions that enforce high-order topology for every band.
Finally, we note that strong $\mathbb{Z}_{2}$ supertopology cannot arise by the mechanism presented in our work. This is because the inversion eigenvalues at the $\Gamma$ point are not constrained by symmetries – they always depend on the particular irrep and are therefore dependent on the orbital character of the bands. However, in higher-dimensional systems our approach can be used to derive symmetry-enforced $\mathbb{Z}_{2}$ topologies in 3D subplanes of the higher-dimensional BZ.